OpenStudy (anonymous):

I can just plug them in and if they work, its right.. right? -x-4y=18 2x-4y=24 A.) (4,-5) B.) (-3,6) C.) (2,-5) D.) No Solution

OpenStudy (anonymous):

i need someone too explain it too me step by step

OpenStudy (anonymous):

just try to substitute the value for y in the main equation 3x+3(-2)=-2 3x-6=-2 3x=4 x=4/3 and you already have the value for y which is -2 then your order pair would be (x,y) which is (4/3,-2)

OpenStudy (anonymous):

oh danget i meant y to eqaul -x ,but i understand same concept right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

k bro thnx

OpenStudy (anonymous):

K now im stuck. 3x+3(-x)=-2 3x-3x=-2 0?=-2 x=

OpenStudy (anonymous):

0 is not equal to -2 then it has no solution

OpenStudy (anonymous):

oh alright thnx